Suppose the market demand and market supply for venison (deer meat) are given by the equations Qd = 38-3P and Qs = P+2.

i. Solve the equilibrium price and quantity.
iii. Government officials fear that too much venison is being consumed at the market equilibrium, leading to concerns about population of deer. In response the government imposes a tax of RM2 on the producers of venison. Solve a new equilibrium after tax.
iii. Plot a new graph to show the movement of equilibrium price and quantity before and after tax.
iv. Calculate consumer and producer burden. Who will absorb most of the tax?
1
Expert's answer
2021-10-08T09:33:53-0400

Solution:

i.). At equilibrium: Qd = Qs

38 – 3P = P + 2

38 – 2 = P + 3P

36 = 4P

P = 9

Equilibrium Price = 9

Substitute in the demand function to derive Quantity:

Qd = 38 – 3P

Qd = 38 – 3(9)

Qd = 38 – 27 = 11

Equilibrium quantity = 11

 

ii.). Rewrite the demand and supply equations as follows:

Demand function: P = "\frac{38}{3} - \frac{Q}{3}"

Supply function: P = Q – 2

New supply function after addition of RM2 tax: P = Q – 2 + 2

Set equal the demand and supply functions:

"\frac{38}{3} - \frac{Q}{3}" = Q – 2 + 2

Q = 9.5

New equilibrium Quantity = 9.5

Substitute to derive the price:

P = "\frac{38}{3} - \frac{9.5}{3}" = 9.5

Equilibrium Price = 9.5

 

iii.). The new graph showing the movement of equilibrium price and quantity before and after tax is as follows:

 

 


iv.). Consumer tax burden = (9.5 – 9) x 9.5 = 0.5 x 9 = 4.75

Consumer tax burden = RM4.75

 

Producer tax burden = (2 – 0.5) x 9.5 = 1.5 x 9.5 = 13.5

Producer tax burden = RM13.5

 

The producer will absorb most of the tax burden.